OpenStudy (anonymous):

how do I simiplify tan(x)^2=sin(x)^2/cos(x)^2 using the expression sin(x)^2+cos(x)^2=1?

OpenStudy (anonymous):

You can solve for sin(x)^2 in the second equation. Then substitute into the first equation.

OpenStudy (anonymous):

i dont get it :(

OpenStudy (anonymous):

sin(x)^2+cos(x)^2=1

OpenStudy (anonymous):

solve for sin(x)^2

OpenStudy (anonymous):

That means get sin(x)^2 by itself on one side of the equation.

OpenStudy (anonymous):

Isolate it!

OpenStudy (anonymous):

i got 1-cos(x)^2

OpenStudy (anonymous):

correct. now substitute what you got for sin(x)^2 in the first equation

OpenStudy (anonymous):

substitute means replace sin(x)^2 with what you got.

OpenStudy (anonymous):

tan(x)^2=sin(x)^2/cos(x)^2=(1-cos(x)^2)/cos(x)^2

OpenStudy (anonymous):

now simplify the right hand side.

OpenStudy (anonymous):

(1-cos(x)^2)/cos(x)^2 = (1/cos(x)^2) - 1

OpenStudy (anonymous):

simplify further to get (1/cos(x)^2) - 1 = sec(x)^2 - 1

OpenStudy (anonymous):

we now have tan(x)^2 = sec(x)^2 -1 add 1 to both sides of the equation to get the identity we seek.

mathslover (mathslover):

Good going cebroski. But confirm in between whether the user is getting your points or not!

OpenStudy (anonymous):

thank you mathslover

OpenStudy (anonymous):

subodha, how are you doing?

OpenStudy (anonymous):

i dont get how u got 1-cos(x)^2/cos(x)^2 to sec(x)^2-1 :(

OpenStudy (anonymous):

Great question! We can distribute the denominator and split the fraction into two fractions.

OpenStudy (anonymous):

A simpler example of this idea is (a - b)/c = a/c - b/c Or, (2 - 1)/3 = 2/3 - 1/3

OpenStudy (anonymous):

think of cos(x)^2 as a single object or thing. In your mind, replace it with a symbol, such as an X. 1-cos(x)^2/cos(x)^2 = (1 - X)/X = 1/X - X/X

OpenStudy (anonymous):

1-cos(x)^2/cos(x)^2 should have parentheses around the numerator.

OpenStudy (anonymous):

(1-cos(x)^2)/cos(x)^2

OpenStudy (anonymous):

how are you doing @subodha ?

OpenStudy (anonymous):

i think i get it thank you :)

OpenStudy (anonymous):

what is your final answer?

OpenStudy (anonymous):

did you end up getting an identity with tangent and secant in it?

OpenStudy (anonymous):

yeah is it becuse 1/cos(x)^2 -1 is equal to sec(x)^2-1

OpenStudy (anonymous):

i get how u got 1-cos(x)^2/cos(x)^2 to sec(x)^2-1

OpenStudy (anonymous):

super. If you add 1 to both sides of the equation, you will get another common pythagorean trigonometric identity.

OpenStudy (anonymous):

oh okay thanks for helping <3